Calculate the potential of a zinc electrode immersed in 0.0200 M in NaOH and saturated with
Zn(OH)2 (E0=Zn2+ / Zn: -0.763 V, Ksp: 3×10-16)
1
Expert's answer
2020-05-24T02:37:45-0400

Zn(OH)2 = Zn2+ + 2OH-

[Zn2+] = Ksp / [OH-]2 = 3x10-16 / (0.0200)2 = 7.5x10-13 M

Zn2+ + 2e = Zn

E = -0.763 - 0.0592/2 x log(1/7.5x10-13) = -1.09 V